Fermi National Laboratory
 

Operation Summary for June 9, 2010 - June 11, 2010

 

Wednesday June 9

The day shift began with Operations monitoring the Recycler stash (270.9E10), with the Antiproton Source (Pbar) stacking (24.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and the Meson FTBF T-977 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store.

During the day shift, Tevatron (TeV) personnel conducted studies until 12:25 PM. Operations put the TeV into HEP shot setup and established store 7882 at 1:45 PM with an average initial luminosity (AIL) of 338E30. Operators prepared the MTest 6 area for T-977 access and detector work; they resumed taking beam at 4 PM.

During the evening shift, Operations disabled the T-977 experiment at 6 PM. The I- Source started to degrade; operators increased its source pressure.

Thursday June 10

The midnight shift began with Operations monitoring store 7882 (91.66E30) and the stash (261.9E10), with Pbar stacking (19.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7882 at 6:05 AM. Operators prepared the CDF and D0 collision halls for accesses at 6:20 AM. At 7:10 AM, Operations reported that a Booster RF station (BRF10) had tripped off numerous times. Operations tuned the Booster RF to operate without the station; further, Operations reduced beam intensities for stacking, NuMI, and MiniBooNE due to the missing RF station.

During the day shift, Booster personnel began their access early due to problems with a Booster RF station (BRF10); they replaced its power amplifier. Booster came back online at 11 AM, with beam to Pbar and MiniBooNE quickly following. NuMI horn power supply work occurred during the Booster access. CDF and D0 came out of access at 11:55 AM. NuMI beam operations resumed at 12:15 PM. Operations put the TeV into shot setup at 2:25 PM and established store 7885 at 3:40 PM with an AIL of 369E30. Operations resumed sending beam to T-977.

During the evening shift, Pbar personnel conducted parasitic beam studies. An instrumentation problem at the Cryo Test Facility required an investigation from cryo system personnel.

Friday June 11

The midnight shift began with Operations monitoring store 7885 (115.1E30) and the stash (398.3E10), with Pbar stacking (24.1E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported that a TeV RF station’s (TRF4) voltage had drooped briefly, increasing the amount of beam in the abort gap. Operators cleared the abort gap and CDF resumed store data acquisition. Operations terminated store 7885 at 6:20 AM and established store 7886 at 7:55 AM with an AIL of 354E30.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/09/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory